# Jet launching from the Kerr black hole magnetosphere: An electrogeodesic approach

### v1

@@ Introduction
There comes then
$A^{(0,\beta)}_{\mu}\textrm{d}x^{\mu}$ $=\beta\chi_{\mu}\textrm{d}x^{\mu}$
$=-\frac{a{\mathcal{B}}Mr\sin^{2}{\theta}}{\Sigma}\,\textrm{d}t+\frac{{\mathcal{B}}\left(r^{4}+a^{2}\Sigma+r^{2}a^{2}\cos^{2}{\theta}+2a^{2}Mr\sin^{2}{\theta}\right)}{2\Sigma}\sin^{2}{\theta}\,\textrm{d}\varphi,$ (2.33)
@@
$A^{\text{W}}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}+\beta\chi_{\mu}\textrm{d}x^{\mu}$ (2.39)
@@
$J^{\mu}\partial_{\mu}=\frac{2}{\Sigma}\left[-\Delta\cos{\theta}\,\partial_{r}+(r-M)\sin{\theta}\,\partial_{\theta}\right].$ (2.43)
@@
Consider a metric which admits a tensor satisfying
$\nabla_{\mu}H_{\nu\alpha}=2g_{\mu[\nu}\xi_{\alpha]},\qquad\nabla_{[\mu}H_{\nu\alpha]}=0,\qquad\xi_{\alpha}=\frac{1}{3}\nabla_{\mu}H^{\mu}{}_{\alpha}$ (2.45)
@@
$Y_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}H^{\rho\sigma},\qquad\nabla_{(\mu}Y_{\nu)\alpha}=0$ (2.46)
@@
$K^{\mu\nu}=Y^{\mu\alpha}Y_{\alpha}{}^{\nu},\qquad\nabla_{(\alpha}K_{\mu\nu)}=0$ (2.47)
which thus provides the Killing tensor for the underlying geometry. Finally, using the principal tensor and the Killing tensor, one can further construct two vectors such that
$\xi^{\mu}\partial_{\mu}$ $=\frac{1}{3}\nabla_{\alpha}H^{\alpha\mu}\partial_{\mu},\qquad\text{such that }\qquad\nabla_{(\mu}\xi_{\nu)}=0$ (2.48)
$\chi^{\mu}\partial_{\mu}$ $=-K^{\mu}{}_{\alpha}\xi^{\alpha}\partial_{\mu},\qquad\text{such that }\qquad\nabla_{(\mu}\chi_{\nu)}=0$ (2.49)
@@
One can easily check that, using the definitions (2.48)–(2.49), one recovers the two commuting Killing vectors
$\xi^{\mu}\partial_{\mu}=\partial_{t}\qquad\chi^{\mu}\partial_{\mu}=\partial_{\varphi}$ (2.52)
Having review these hidden Killing objects of the Kerr geometry, we can now discuss two alternatives ways to construct exact solutions of the Maxwell equations on the Kerr geometry.
@@
Consider the principal tensor $H$ of the Kerr geometry given by (2.50). It provides an exact solution of the non-vacuum Maxwell equation known as the Killing-Maxwell (KM) solution. To see this, notice first that by definition, it is a closed form so that it satisfies the Bianchi identify $\textrm{d}H=0$. Let us introduce the associated Faraday tensor
$F^{\text{KM}}=\frac{1}{3}H_{\mu\nu}\textrm{d}x^{\mu}\wedge\textrm{d}x^{\nu}=\frac{1}{3}\left(re^{0}\wedge e^{1}+a\cos{\theta}\;e^{2}\wedge e^{3}\right)$ (2.53)
which can be derived from the gauge field
@@
Then, using the last equation of (2.45), it can be recast as
$\nabla_{\mu}F_{\text{KM}}^{\mu\nu}=J^{\nu}\qquad\text{with}\qquad J^{\mu}\partial_{\mu}=\partial_{t}$ (2.55)
such that the Maxwell-Gauss equation is satisfied for an electric current given precisely by the time-like Killing vector $\xi^{\mu}\partial_{\mu}=\partial_{t}$. It is remarkable that the fundamental Killing object of the Killing tower provides without any further condition an exact non-vacuum solution of the test Maxwell equation. This system was studied by Carter in [73] where he showed that even more remarkably, it provides one example of a magnetosphere for the Kerr black hole which preserves the existence of the Carter constant for the electrogeodesic motion. The same construction was later studied by Krtous in [61] to build Maxwell solutions on higher dimensional rotating black holes.
@@
Another approach to construct vacuum Maxwell solution in a metric possessing a KY tensor was introduced by Penrose in [63]. Consider the following antisymmetric rank-$2$ tensor
$F_{\mu\nu}=\alpha\;R_{\mu\nu\alpha\beta}Y^{\alpha\beta}$ (2.56)
@@
$\nabla_{\mu}F^{\mu\nu}$ $=R^{\rho\sigma\mu\nu}\nabla_{\mu}Y_{\rho\sigma}+Y^{\rho\sigma}\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}$ (2.57)
$=-R^{\nu\mu\rho\sigma}\nabla_{[\mu}Y_{\rho\sigma]}+Y^{\rho\sigma}\cancel{\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}}$ (2.58)
$=-R^{\nu[\mu\rho\sigma]}\nabla_{\mu}Y_{\rho\sigma}$ (2.59)
$=0$ (2.60)
Here, we have used the cyclic properties of the Riemann, i.e. $R_{\mu[\nu\rho\sigma]}=0$, the identify $\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}=0$ which holds only in vacuum, i.e. when $R_{\alpha\beta}=0$ and the definition of the KY tensor given by
$\nabla_{(\mu}Y_{\nu)\alpha}=0\qquad\nabla_{\mu}Y_{\nu\alpha}=\nabla_{[\mu}Y_{\nu\alpha]}=0\qquad\nabla_{\mu}Y^{\mu}{}_{\alpha}=0$ (2.61)
Taking the Hodge dual, one finally obtains that
$\nabla_{\mu}F^{\mu\nu}=0\qquad\nabla_{\mu}\tilde{F}^{\mu\nu}=0$ (2.62)
where $\tilde{F}_{\mu\nu}=\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ is the dual tensor. Interestingly, applying this to the Kerr geometry and using $\alpha=-{\mathcal{P}}/(2M)$, one obtains [^2] See Notebook 1 in App. A for the computation.
@@
which reproduces the solution (2.24) discussed in the previous section. Therefore, the Penrose current of the Kerr spacetime coincides with the dual solution of the Wald construction based solely on the time-like Killing vector $\xi^{\mu}\partial_{\mu}$.
@@
This covers in particular the case of the test Kerr magnetic monopole which will be the central magnetosphere model we will study.
@@
A set of conditions can be derived for the geodesic motion of neutral particles on the Kerr geometry. They descend from the basic requirement that the $4$-velocity $u^{\mu}\partial_{\mu}$ is a time-like vector. They can be expressed in terms of the energy and the angular momentum of the neutral massive particle. As they have important consequences for the geodesic motion, our goal here is to discuss how these constraints are modified for electrogeodesics.
To start with, let us first define the conserved energy $E$ and conserved angular momentum $L$ of the charged particle. As usual, the Kerr geometry being stationary and axisymmetric, one can use its timelike Killing vector $\xi^{\mu}\partial_{\mu}=\partial_{t}$ and the azimuthal Killing vector $\chi^{\mu}\partial_{\mu}=\partial_{\varphi}$ to build $E$ and $L$. Using the canonical momenta (3.4), one obtains
@@
They provide the corrected notion of energy and momenta in the presence of the Maxwell field.
Let us now review the different conditions on the electrogeodesic.
@@
An important property of the geodesic motion on the Kerr spacetime is exhibit a set of four constants of motion which are in involution. Three of them are the energy $E$, the angular momentum $L$ which descend from the time and azimuthal isometries of the Kerr geometry. The third one is the mass $m$ of the particle. However, the fourth constant of motion was found by Carter in [54] and descends from the existence of the hidden symmetry of the Kerr geometry giving rise to the Killing tower reviewed in Section 2.2.1. Concretely, for a neutral test particle with momenta $p_{\mu}$, the Killing tensor (2.47) gives rise to a constant ${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu}$ along the geodesic flow, i.e. $D{\mathcal{K}}=0$, known as the Carter constant.
@@
$=\frac{1}{6}\nabla^{(\alpha}K^{\mu\nu)}p_{\alpha}p_{\mu}p_{\nu}+F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu}$ (3.26)
@@
$\{{\mathcal{K}},H\}$ $=F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu}$ (3.27)
@@
$\{{\mathcal{K}},H\}=0\qquad\text{if}\qquad F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0$ (3.28)
@@
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\alpha}Y^{\beta\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\nu}Y^{\beta\mu})Y^{\alpha}_{\beta}$ (3.29)
@@
$S^{\alpha\beta}=F_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0$ (3.30)
@@
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\beta}Y^{\alpha\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\beta}Y^{\nu\mu})Y^{\alpha}_{\beta}=Y^{\alpha\beta}Y^{\nu\mu}\left(F_{\beta\mu}+F_{\mu\beta}\right)=0$ (3.31)
Therefore, the condition (3.30) is sufficient to ensures that $F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0$, as first noticed in [74]. This condition selects the models of Kerr magnetosphere which preserve the Carter constant.
We can now check which solution for the Kerr magnetosphere satisfies this peculiar condition. Let us examine each magnetosphere models discussed in the previous section.
@@
    $F_{\alpha\beta}=H_{\alpha\beta},\qquad S^{\alpha\beta}=H_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0$ (3.32)
@@
  • The Kerr magnetic monopole (and Penrose current): Consider now the magnetosphere built from the Penrose current (or the dual time-like Killing vector solution). One has
    $F_{\mu\nu}=R_{\mu\nu\rho\sigma}Y^{\rho\sigma}\qquad S^{\alpha\beta}=R_{\mu}{}^{[\alpha}{}_{\rho\sigma}Y^{\rho\sigma}Y^{\beta]\mu}=0$ (3.33)
@@
  • Wald solution: Consider now the Wald solution. One finds [^4] See Notebook 1 in App. A for the computation. that $S^{\alpha\beta}\neq 0$ in this case: the 2-form $\underline{S}=S_{\alpha\beta}\;\textrm{d}x^{\alpha}\wedge\textrm{d}x^{\beta}$ metric-dual to $S^{\alpha\beta}$ is
@@
Having reviewed the conditions for a magnetosphere to preserve the Carter constant, and demonstrated that the Kerr monopole solution belongs to this very special class, we naturally turn towards writing the first order electrogeodesic equations.
@@
The main goal of this section is to review the different steps to separate of the electrogeodesic equations within the Kerr test monopole magnetosphere. The first order electrogeodesic equations we shall derive are a subset of the electrogeodesic equations of the Kerr-Newman black hole [55] where one set the electric charge to zero and keep the magnetic charge arbitrary.
@@
In order to separate the radial equation, we first consider first the Carter constant given by
${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu}$ (3.35)
where $K_{\mu\nu}$ is the Killing tensor of the Kerr geometry which can be decompose as
$K_{\alpha\beta}=(r^{2}+a^{2})(k_{\alpha}\ell_{\beta}+\ell_{\alpha}k_{\beta})+r^{2}g_{\alpha\beta}$ (3.36)
Here $k^{\alpha}\partial_{\alpha}$ and $\ell^{\alpha}\partial_{\alpha}$ are the null vectors corresponding to the principal null directions of Kerr geometry. They are given explicitly by
@@
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})(E+qA_{t})-a(L-qA_{\varphi})\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.39)
$=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL+{q((r^{2}+a^{2})A_{t}+aA_{\varphi})}\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.40)
such that one obtains a non-trivial contribution from the magnetosphere encoded in the last term. Interestingly, focusing on the Kerr monopole, one finds that this term vanishes,
$(r^{2}+a^{2})A_{t}+aA_{\varphi}=0$ (3.41)
so that the Carter constant takes the same form as for geodesic
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.42)
Since this constant depends only on the radial component $p^{r}$, it allows one to algebraically deduce the first order radial equation which reads
$\Sigma\;p^{r}=\Sigma m\frac{\textrm{d}r}{\textrm{d}\lambda}=\sqrt{R(r)}\qquad\text{with}\qquad R(r)=\left[(r^{2}+a^{2})E-aL\right]^{2}-\Delta(r^{2}m^{2}+{\mathcal{K}})$ (3.43)
This equation reproduces the one of the geodesic case, but one has to keep in mind that now, it involves a corrected notion of the energy and angular momentum. Notably, the l.h.s only involves functions of the radius $r$.
Now, we can obtain the first order equation for the polar motion using both the Carter constant and the mass relation $m^{2}=-g_{\mu\nu}p^{\mu}p^{\nu}$. Writing this explicitly, one finds that
$m^{2}=-g^{\mu\nu}\left(P_{\mu}-qA_{\mu}\right)\left(P_{\nu}-qA_{\mu}\right)=-g^{\mu\nu}P_{\mu}P_{\nu}+2qg^{\mu\nu}P_{\mu}A_{\nu}-q^{2}g^{\mu\nu}A_{\mu}A_{\nu}$ (3.44)
Since $A_{r}=A_{\theta}=0$ so that $P_{r}=p_{r}$ and $P_{\theta}=p_{\theta}$, the first term will reproduce the standard term for the geodesic. The challenge lies in the two last terms which need to provide separable form. Concretely, they are given by
@@
$g^{\mu\nu}A_{\mu}A_{\nu}$ $=\frac{{\mathcal{P}}^{2}}{\Sigma}\frac{\cos^{2}{\theta}}{\sin^{2}{\theta}}$ (3.46)
@@
$\qquad-\frac{\Sigma}{\Delta}(p^{r})^{2}-\Sigma(p^{\theta})^{2}$ (3.47)
@@
with the potential given by
$\Theta(\theta)={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos\theta}{\sin{\theta}}-aE\sin{\theta}\right)^{2}$ (3.50)
The polar motion follows the same equation as for the geodesic which also involves only functions of the polar angle $\theta$. In conclusion, the property (3.30) provides us with a Carter-like constant which, together with the mass, can be used to derive the first order equations for the radial and polar motion. Both equations are separable and remain unchanged w.r.t to the geodesic motion. This already implies that the different radial and polar equilibrium positions for the charged particles will remain the same as the one of their neutral counterpart. We shall come back on this point in the following.
@@
Let us start with the equations for $\dot{t}$ and $\dot{\varphi}$. The expressions for the energy and the angular momentum can be written as
$\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right)$ $=\left(\begin{array}[]{cccc}g_{tt}&g_{t\varphi}\\ g_{\varphi t}&g_{\varphi\varphi}\end{array}\right)\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right)$ (3.57)
@@
$\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right)=\frac{1}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}\left(\begin{array}[]{cccc}g_{\varphi\varphi}&-g_{t\varphi}\\ -g_{\varphi t}&g_{tt}\end{array}\right)\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right)$ (3.58)

### v2

@@ Introduction
Since $A^{(0,\beta)}_{\mu}=\beta\chi_{\mu}$, there comes
$A^{(0,\beta)}=-\frac{a{\mathcal{B}}Mr\sin^{2}{\theta}}{\Sigma}\,\textrm{d}t+\frac{{\mathcal{B}}\left(r^{4}+a^{2}\Sigma+r^{2}a^{2}\cos^{2}{\theta}+2a^{2}Mr\sin^{2}{\theta}\right)}{2\Sigma}\sin^{2}{\theta}\,\textrm{d}\varphi,$ (2.32)
@@
$A^{\text{W}}_{\mu}=\alpha\xi_{\mu}+\beta\chi_{\mu},$ (2.38)
@@
$J=\frac{2}{\Sigma}\left[-\Delta\cos{\theta}\,\partial_{r}+(r-M)\sin{\theta}\,\partial_{\theta}\right].$ (2.42)
@@
Consider a metric which admits a rank-2 tensor $H$ satisfying
$\nabla_{\mu}H_{\nu\alpha}=2g_{\mu[\nu}\xi_{\alpha]},\qquad\nabla_{[\mu}H_{\nu\alpha]}=0,\qquad\xi_{\alpha}=\frac{1}{3}\nabla_{\mu}H^{\mu}{}_{\alpha}$. (2.44)
@@
$Y_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}H^{\rho\sigma},\qquad\nabla_{(\mu}Y_{\nu)\alpha}=0,$ (2.45)
@@
$K^{\mu\nu}=Y^{\mu\alpha}Y_{\alpha}{}^{\nu},\qquad\nabla_{(\alpha}K_{\mu\nu)}=0,$ (2.46)
which thus provides the Killing tensor for the underlying geometry. Finally, using the principal tensor and the Killing tensor, one can further construct two vectors $\xi$ and $\chi$ such that
$\xi^{\mu}$ $=\frac{1}{3}\nabla_{\alpha}H^{\alpha\mu},\qquad\text{such that }\qquad\nabla_{(\mu}\xi_{\nu)}=0$ (2.47)
$\chi^{\mu}$ $=-K^{\mu}{}_{\alpha}\xi^{\alpha},\qquad\text{such that }\qquad\nabla_{(\mu}\chi_{\nu)}=0.$ (2.48)
@@
One can easily check that, using the definitions (2.47)–(2.48), one recovers the two commuting Killing vectors (2.10). Having reviewed these hidden Killing objects of the Kerr geometry, we can now discuss two alternatives ways to construct exact solutions of the Maxwell equations on the Kerr geometry.
@@
Consider the principal tensor $H$ of the Kerr geometry given by (2.49). It provides an exact solution of the non-vacuum Maxwell equation known as the Killing-Maxwell (KM) solution [75]. To see this, notice first that by definition, it is a closed form: $\textrm{d}H=0$. Let us introduce the associated KM Faraday tensor by
$F^{\text{KM}}=\frac{1}{3}H=\frac{1}{3}\left(re^{0}\wedge e^{1}+a\cos{\theta}\;e^{2}\wedge e^{3}\right).$ (2.51)
It can be derived from the 4-potential
@@
Then, using the last equation of (2.44), we get
$\nabla_{\nu}F_{\text{KM}}^{\mu\nu}=4\pi J^{\mu}\qquad\text{with}\qquad J=\frac{1}{4\pi}\partial_{t}.$ (2.53)
Hence $F^{\text{KM}}$ fulfills the Maxwell equations with an electric 4-current given by the asymptotically timelike Killing vector $\xi=\partial_{t}$, up to a $4\pi$ factor. It is remarkable that the fundamental Killing object of the Killing tower provides without any further condition an exact non-vacuum solution of the test Maxwell equation. This system was studied by Carter [75], who showed that even more remarkably, it provides one example of a magnetosphere for the Kerr black hole which preserves the existence of the Carter constant for the electrogeodesic motion. The same construction was later studied by Krtous in [63] to build Maxwell solutions on higher dimensional rotating black holes.
@@
Another approach to construct vacuum Maxwell solutions in a metric possessing a KY tensor $Y$ was introduced by Penrose in [65]. Consider the following antisymmetric rank-$2$ tensor
$F_{\mu\nu}=\alpha\;R_{\mu\nu\alpha\beta}Y^{\alpha\beta}.$ (2.54)
@@
$\nabla_{\mu}F^{\mu\nu}$ $=R^{\rho\sigma\mu\nu}\nabla_{\mu}Y_{\rho\sigma}+Y^{\rho\sigma}\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}$
$=-R^{\nu\mu\rho\sigma}\nabla_{[\mu}Y_{\rho\sigma]}+Y^{\rho\sigma}\cancel{\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}}$
$=-R^{\nu[\mu\rho\sigma]}\nabla_{\mu}Y_{\rho\sigma}$
$=0.$ (2.55)
Here, we have used the cyclic properties of the Riemann tensor, i.e. $R_{\mu[\nu\rho\sigma]}=0$, the identity $\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}=0$, which holds only in vacuum (i.e. when $R_{\alpha\beta}=0$), and the definition of the KY tensor given by
$\nabla_{(\mu}Y_{\nu)\alpha}=0\qquad\nabla_{\mu}Y_{\nu\alpha}=\nabla_{[\mu}Y_{\nu\alpha]}=0\qquad\nabla_{\mu}Y^{\mu}{}_{\alpha}=0.$ (2.56)
Taking the Hodge dual, one finally obtains
$\nabla_{\mu}F^{\mu\nu}=0\quad\mbox{and}\quad\nabla_{\mu}*F^{\mu\nu}=0,$ (2.57)
where $F_{\mu\nu}:=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ is the dual tensor. Interestingly, applying this to the Kerr geometry and using $\alpha=-{\mathcal{P}}/(2M)$ in (2.54), one obtains [^2] See Notebook 1 in App. A for the computation.
@@
which reproduces the solution (2.23) discussed in the previous section. Therefore, the Penrose current of the Kerr spacetime coincides with the dual solution of the Wald construction based solely on the stationary Killing vector $\xi$. This concludes the review on the construction of exact Maxwell solutions from the underlying symmetries. Let us now discuss how to study the motion of charges particles in such magnetospheres.
@@
w.r.t the Boyer-Lindquist coordinates. This covers in particular the case of the test Kerr magnetic monopole which will be the central magnetosphere model we will study.
@@
A set of conditions can be derived for the geodesic motion of neutral particles on the Kerr geometry. They descend from the basic requirement that the $4$-velocity $u^{\mu}$ is a timelike vector. They can be expressed in terms of the energy and the angular momentum of the neutral massive particle. As they have important consequences for the geodesic motion, our first goal here is to discuss how these constraints are modified for electrogeodesics.
To start with, let us first define the conserved energy $E$ and conserved angular momentum $L$ of the charged particle. As usual, the Kerr geometry being stationary and axisymmetric, one can use its time translation Killing vector $\xi=\partial_{t}$ and its azimuthal Killing vector $\chi=\partial_{\varphi}$ to build $E$ and $L$. Using the canonical momenta (3.4), one obtains
@@
$E$ and $L$ provide the corrected notion of energy and momenta in the presence of the Maxwell field.
Let us now review the different conditions on the electrogeodesics.
@@
An important property of the geodesic motion on the Kerr spacetime is that it exhibits a set of four constants of motion which are in involution. Two of them are the energy $E$, the angular momentum $L$ which descend from the time and azimuthal isometries of the Kerr geometry. The third one is the mass $m$ of the particle. However, the fourth constant of motion was found by Carter in [55] and descends from the existence of the hidden symmetry of the Kerr geometry giving rise to the Killing tower reviewed in Section 2.2.1. Concretely, for a neutral test particle with momenta $p_{\mu}$, the Killing tensor (2.46) gives rise to a constant ${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu}$ along the geodesic flow, i.e. $D{\mathcal{K}}=0$, known as the Carter constant, where ${\mathcal{D}}=u^{\mu}\nabla_{\mu}$ is the derivative along the geodesic flow.
@@
$=\frac{1}{6}\nabla^{(\alpha}K^{\mu\nu)}p_{\alpha}p_{\mu}p_{\nu}+F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu},$ (3.26)
@@
$\{{\mathcal{K}},H\}$ $=F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu},$ (3.27)
@@
$\{{\mathcal{K}},H\}=0\qquad\text{if}\qquad F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0.$ (3.28)
@@
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\alpha}Y^{\beta\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\nu}Y^{\beta\mu})Y^{\alpha}_{\beta}.$ (3.29)
@@
$S^{\alpha\beta}=F_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0.$ (3.30)
@@
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\beta}Y^{\alpha\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\beta}Y^{\nu\mu})Y^{\alpha}_{\beta}=Y^{\alpha\beta}Y^{\nu\mu}\left(F_{\beta\mu}+F_{\mu\beta}\right)=0.$ (3.31)
Therefore, the condition (3.30) is sufficient to ensures that $F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0$, as first noticed in [76]. This condition selects the models of Kerr magnetosphere that preserve the Carter constant. Notice that while we demand here that the Carter constant be preserved, it is in principle possible that for some magnetosphere models, the electrogeodesic is separable for a corrected version of the Carter constant. The condition (3.30) only applies to the former case.
We can now check which solution for the Kerr magnetosphere satisfies this peculiar condition. Let us examine each magnetosphere model discussed in the previous section.
@@
    $F_{\alpha\beta}=H_{\alpha\beta},\qquad S^{\alpha\beta}=H_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0.$ (3.32)
@@
  • The Kerr magnetic monopole (and Penrose current): Consider now the magnetosphere built from the Penrose current (or the dual timelike Killing vector solution). One has
    $F_{\mu\nu}=R_{\mu\nu\rho\sigma}Y^{\rho\sigma},\qquad S^{\alpha\beta}=R_{\mu}{}^{[\alpha}{}_{\rho\sigma}Y^{\rho\sigma}Y^{\beta]\mu}=0.$ (3.33)
@@
  • Wald solution: Consider now the Wald solution. One finds [^4] See Notebook 1 in App. A for the computation. that $S^{\alpha\beta}\neq 0$ in this case: the 2-form $S_{\alpha\beta}$ metric-dual to $S^{\alpha\beta}$ is
@@
    This explains why the electrogeodesic motion is not separable on the Wald solution.
Having reviewed the conditions for a magnetosphere to preserve the Carter constant, and checked that the Kerr monopole solution belongs to this very special class, we naturally turn towards writing the first order electrogeodesic equations.
@@
The main goal of this section is to review the different steps to separate the electrogeodesic equations within the Kerr test monopole magnetosphere. The first order electrogeodesic equations we shall derive are a subset of the electrogeodesic equations of the Kerr-Newman black hole [56] where the electric charge is set to zero and the magnetic charge is kept arbitrary.
@@
In order to separate the radial equation, we first consider the Carter constant given by
${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu},$ (3.35)
where $K_{\mu\nu}$ is the Killing tensor of the Kerr geometry which can be decomposed as
$K_{\alpha\beta}=(r^{2}+a^{2})(k_{\alpha}\ell_{\beta}+\ell_{\alpha}k_{\beta})+r^{2}g_{\alpha\beta}.$ (3.36)
Here $k^{\alpha}$ and $\ell^{\alpha}$ are the null vectors corresponding to the principal null directions of the Kerr geometry; they are given explicitly by
@@
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})(E+qA_{t})-a(L-qA_{\varphi})\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$
$=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL+{q((r^{2}+a^{2})A_{t}+aA_{\varphi})}\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2},$ (3.39)
such that one obtains a non-trivial contribution from the magnetosphere encoded in the last term. Interestingly, focusing on the Kerr monopole, one finds that this term vanishes:
$(r^{2}+a^{2})A_{t}+aA_{\varphi}=0,$ (3.40)
so that the Carter constant takes the same form as for a geodesic:
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}.$ (3.41)
Since this constant depends only on the radial component $p^{r}$, it allows one to algebraically deduce the first order radial equation, which reads
$\Sigma\;p^{r}=\Sigma m\frac{\textrm{d}r}{\textrm{d}\lambda}=\sqrt{R(r)}\qquad\text{with}\qquad R(r)=\left[(r^{2}+a^{2})E-aL\right]^{2}-\Delta(r^{2}m^{2}+{\mathcal{K}}).$ (3.42)
This equation reproduces that of the geodesic case; notably, the l.h.s. only involves functions of the radial coordinate $r$. Nevertheless, one has to keep in mind that now, it involves a corrected notion of the energy and angular momentum.
Now, we can obtain the first order equation for the polar motion using both the Carter constant and the mass relation $m^{2}=-g_{\mu\nu}p^{\mu}p^{\nu}$. Writing this explicitly, we find
$m^{2}=-g^{\mu\nu}\left(P_{\mu}-qA_{\mu}\right)\left(P_{\nu}-qA_{\mu}\right)=-g^{\mu\nu}P_{\mu}P_{\nu}+2qg^{\mu\nu}P_{\mu}A_{\nu}-q^{2}g^{\mu\nu}A_{\mu}A_{\nu}.$ (3.43)
Since $A_{r}=A_{\theta}=0$ so that $P_{r}=p_{r}$ and $P_{\theta}=p_{\theta}$, the first term will reproduce the standard term for the geodesic. The challenge lies in the last two terms which need to provide separable form. Concretely, they are given by
@@
$g^{\mu\nu}A_{\mu}A_{\nu}$ $=\frac{{\mathcal{P}}^{2}}{\Sigma}\frac{\cos^{2}{\theta}}{\sin^{2}{\theta}}.$ (3.45)
@@
$\qquad-\frac{\Sigma}{\Delta}(p^{r})^{2}-\Sigma(p^{\theta})^{2}.$ (3.46)
@@
with the latitudinal potential $\Theta(\theta)$ defined by
$\Theta(\theta)={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos\theta}{\sin{\theta}}-aE\sin{\theta}\right)^{2}.$ (3.49)
Thus, contrary to the radial motion, the polar electrogeodesic motion is corrected w.r.t the geodesic motion. However, it also involves only functions of the polar angle $\theta$. In conclusion, the property (3.30) provides us with a Carter-like constant which, together with the mass, can be used to derive the first order equations for the radial and polar motion. Both equations are separable. Since the radial equation remains the same, this already implies that the different radial equilibrium positions for the charged particles will remain the same as the one of their neutral counterpart. On the contrary, the polar equilibrium positions will be modified. We shall come back on this point in the following.
@@
Let us now turn to the equations for $\dot{t}$ and $\dot{\varphi}$. The expressions for the energy and the angular momentum can be written as
$\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right)$ $=\left(\begin{array}[]{cccc}g_{tt}&g_{t\varphi}\\ g_{\varphi t}&g_{\varphi\varphi}\end{array}\right)\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right).$ (3.56)
@@
$\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right)=\frac{1}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}\left(\begin{array}[]{cccc}g_{\varphi\varphi}&-g_{t\varphi}\\ -g_{\varphi t}&g_{tt}\end{array}\right)\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right),$ (3.57)